Skip to end of metadata
Go to start of metadata

First day for the entire crew in the field.  Everyone came so that different perspectives could be shared about tower operations.  All this was done at rsw04, which is now ready for power.  We powered it up with a battery for about 30min to check out everything.  (ARL now is using DTU's portable generator.)

Errors found and fixed at rsw04:

  • eolupdate had reverted to the wrong serial ports for the sonic and licor
  • licor DID changed to reflect height of 10m
  • TP01 had been plugged into incorrect mote port (stupid me!)
  • 15m cables were fine to get to the rad and soil motes

Back in the ops center:

  • Finished prep of our next two towers: tse07 and tse08.  
    • Found a bad Tsoil probe (reporting all -0.001)
    • Found that the PTB110 connector inside the EC100 was loose (just like VERTEX s15!)
    • EC100 needed to be changed to binary mode
    • Found that the EC100 had to be grounded both to the DSM and to the CSAT3A head, when using a long serial cable
    • Changed to longer cables and configuration, since tse07 is now a 30m tower (picklist had shown 20m)
    • One CSAT3 (older) seemed bad and another became bad during testing.  We're hoping that it will act better deployed on a tower.
    • One bad port found on tse08's DSM, so METEK was moved to a different port.  I think this is the first bad port we've seen?
  • Updated software on tnw07's 2 DSMs
  • Reset 16 EC100 boxes to binary mode (needed for about half of these boxes – very strange)
  • Prepped Ubiquiti's for tnw09, tnw11, and rne07, which we hope to get to on Wed.

Interacted a bit with both Samortecnica's 3-person crew and ARL's crew (saw Chris just before he left).  Samortecnica was working on rsw05, 06, 07 today.  ARL has the SW energy balance 10m tower up, though not completely connected; one 30m tower up; at least 3 other 30m towers are erect, but not yet telescoped to full height as their crew is adding instrumentation.  So...progress is being made!